Please help me with this problem. The Width of a rectangle is 3 feet shorter than its length. The perimeter is 306 feet. Let x equal the length of the rectangle. The formula is p= 2L+2w. What is the length and the width of the rectangle?
Answer by rothauserc(4718)

w = x - 3
:
306 = 2x + 2(x - 3)
:
306 = 2x + 2x - 6
:
4x = 312
:
x = 78
:
******************************
length of rectangle = 78 feet
width of rectangle = 75 feet
:
check our answer
:
306 = 2 * 78 + 2 * 75
:
306 = 156 + 150
:
306 = 306
:
our answer checks
******************************